# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

### 19 Mar 2026 v1

@@ 4 Regret Guarantees for Nonstationary Satisficing Bandits
In this section, we establish minimax satisficing regret guarantees for general nonstationary satisficing bandits. Sections 4.1 and 4.2 present the corresponding lower and upper bounds, respectively.
@@ 4.1 Lower Bounds
In the stationary setting, the minimax-optimal satisficing regret is constant under realizability (Michel et al. 2022, Feng et al. 2024, Rouyer et al. 2024). It is therefore natural to conjecture that, in a piecewise-stationary environment with $L$ segments, an $\mathcal{O}(L)$ guarantee might be achieved by restarting a stationary algorithm at each change-point (if the change-points can be efficiently detected). The main message of this subsection is that such an $\mathcal{O}(L)$ rate is not achievable in general when change-points are unknown: an additional $\log T$ factor is unavoidable.
A canonical hard class.
Let $\mathcal{E}_{L,\Delta}$ denote the class of two-armed, piecewise-stationary bandit instances with $L$ segments, where rewards are Gaussian with unit variance and the mean vector on each segment $\ell$ is either $(S+\Delta,S-\Delta)$ or $(S-\Delta,S+\Delta)$. In particular, the unique satisficing arm swaps at every change-point.
This class is canonical for our problem. It is the simplest symmetric realizable family that isolates the effect of nonstationarity: there are only two arms, the gap to the threshold is fixed at $\Delta$ on every segment, and each segment contains exactly one satisficing arm and one non-satisficing arm. Thus, the difficulty arises solely from the unknown change-points and the resulting switches in the identity of the satisficing arm. Furthermore, we emphasize that classical pairwise KL change-of-measure arguments do not directly yield the desired lower bound on this class. Indeed, if two instances in $\mathcal{E}_{L,\Delta}$ differ only in the location of one switch, then on the interval where the switch locations disagree. Hence, under unit-variance Gaussian rewards, each round in this disagreement interval contributes $2\Delta^{2}$ to the conditional KL divergence, regardless of which arm is pulled. As a result, the pairwise KL is driven primarily by the length of the disagreement interval, rather than by the number of regretful pulls. This breaks the usual mechanism behind classical KL-based lower bounds (Bubeck et al. 2012, Garivier et al. 2019, Garivier and Moulines 2011). We therefore use our post-interaction-reference framework to derive the sharp satisficing regret lower bound for $\mathcal{E}_{L,\Delta}$.
@@
When $\Delta$ is a constant and $L\ll\Delta^{2}T$, Theorem 1 implies an $\Omega(L\log T)$ lower bound for general nonstationary satisficing bandits. Relative to the naive $\mathcal{O}(L)$ guess, the extra $\log T$ factor reflects the fact that the change-points are unknown: minimizing satisficing regret is at least as hard as localizing these change-points, which already incurs an $\Omega(L\log T)$ cost. The proof, deferred to Section 6.1, proceeds via Approach I in Section 3.2.
The proof of Theorem 1 is restricted to $L\geq 3$ because it relies on the particular instance construction (13), which is tailored to enforce the separation condition (14) and hence the reduction (10). When $L=2$ (i.e., the environment changes only once), it is unclear how to design an instance family admitting a comparable separation condition; see Figure 1. This leaves open whether the $\log T$ factor is also unavoidable when $L=2$. The next theorem closes this gap by establishing the same scaling for all $L\geq 2$, under a slightly stronger condition relating $\Delta$, $T$, and $L$.
separationseparationL=3L=3no separationL=2L=2
Figure 1: Schematic illustration: for $L=3$ one can create a separated local window to enable a clean identification reduction; for $L=2$ the analogous separation fails (motivating the information-budget approach).
@@
Theorem 2 follows Approach II in Section 3.2; details are in Section 6.2. It also offers complementary intuition for the $\Omega(L\log T)$ scaling: regardless of the policy, one always acquires $\Omega(\log T)$ information per segment, since the reference can be used to drive the identification error below a constant. The information-cost bound then shows that obtaining this information necessarily incurs $\Omega(\log T)$ satisficing regret per segment.
@@ 4.2 Upper Bounds
We now present an algorithm whose satisficing regret matches the lower bounds above (up to problem-dependent constants). We begin by defining the windowed statistics used throughout.
Windowed statistics.
For any buffer $\mathcal{B}$ of length $m$, define the dyadic window set $\mathrm{WinSet}(\mathcal{B}):=\{2^{k}:2^{k}\leq m,\ k\in\mathbb{N}\}$. For each $w\in\mathrm{WinSet}(\mathcal{B})$, define the average of the most recent $w$ entries by $\mathrm{AvgLast}(\mathcal{B},w):=\frac{1}{w}\sum_{s=m-w+1}^{m}\mathcal{B}[s].$ Given a confidence radius $\beta(\cdot)$, define the windowed lower and upper confidence bounds $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B}):=\max_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)-\beta(w)\big\}$ and $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})=\min_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)+\beta(w)\big\}.$
@@
Algorithm 1 maintains a leader state. When $\textsf{leader}=\texttt{None}$, it explores until it certifies that an arm is satisficing. When a leader is present, the algorithm exploits this arm while continuously testing whether it remains satisficing. The key difficulty in piecewise-stationary environments is that, after a change-point, stale samples can be misleading. The windowed bounds address this issue: $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})$ takes the most optimistic lower bound over dyadic suffix windows, while $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})$ takes the most pessimistic upper bound. Consequently, if an arm’s current mean exceeds $S$, then eventually $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})\geq S$; conversely, if the current leader becomes non-satisficing, then for a sufficiently recent window we have $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})<S$, triggering a reset and renewed exploration.
We now state the regret guarantee for Algorithm 1.
@@
Under Assumption 1, let $\pi$ be the Algorithm 1. Then,
@@
The logarithmic dependence on $T$ arises from concentration: with $\beta(w)\asymp\sqrt{\log T/w}$, certifying $\mu>S$ (via $\mathrm{LCB}^{\operatorname{win}}\geq S$) or rejecting $\mu<S$ (via $\mathrm{UCB}^{\operatorname{win}}<S$) requires only $w=\mathcal{O}(\log T/\Delta^{2})$ samples, where $\Delta$ is the relevant gap to the threshold. The windowed bounds also prevent frequent resets: after a reset, if the environment remains unchanged, then with high probability a satisficing leader is not demoted and a non-satisficing candidate is not promoted. Thus, within each stationary segment, the algorithm spends only a logarithmic number of rounds (i) exploring until it finds a satisficing leader, or (ii) detecting that the current leader has fallen below $S$; on the remaining rounds it plays a satisficing arm and incurs no additional regret. The full proof of Theorem 3 is deferred to Section 6.3.
5 When Is Constant Satisficing Regret Achievable
It remains unclear which structural conditions permit constant satisficing regret in nonstationary $K$-armed bandits. In the stationary setting, constant satisficing regret is possible because the learner can identify satisficing and non-satisficing arms in finite expected time—independent of TT—by comparing empirical means to the threshold $S$; see (Michel et al. 2022, Appendix A).
A common feature of the hard instances behind the $\Omega(\log T)$ lower bounds in Section 4 is the presence of threshold crossings: an arm may be satisficing early on but later down-cross below $S$ at an unknown time, and the identity of the realizable arm may also switch across segments. Such changes force the learner to repeatedly certify satisficing arms, which incurs a logarithmic cost.
In this section, we identify a simple regime in which this obstacle disappears. Specifically, if (i) there exists an always-realizable arm that remains strictly satisficing throughout, and (ii) no arm ever transitions from satisficing to non-satisficing (i.e., no down-crossings), then constant satisficing regret—independent of both the horizon $T$ and the number of stationary segments LL—is achievable.
Formally, we impose the following assumption.
@@ 5 When Is Constant Satisficing Regret Achievable
Under Assumptions 2 and 3, let $\pi$ be Algorithm 2. Then,
@@
Theorem 4 establishes a constant satisficing regret bound that does not depend on $T$ or $L$, and it recovers the stationary guarantee when $L=1$; see Michel et al. (2022). The key proof idea is that, under Assumptions 2 and 3, the always-realizable arm and any non-satisficing arm can be identified in finite expected time before the latter ever becomes satisficing. The full proof is deferred to Section 6.4.
We emphasize that Assumptions 2 and 3 are sufficient (but not necessary) for constant satisficing regret. Characterizing necessary and sufficient structural conditions remains an interesting direction for future work.
@@ 6.1 Proof of Theorem 1
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
We assume Gaussian rewards of the form $r_{t}(a)=\mu_{t}(a)+\xi_{t}$, where $\xi_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume $L$ is odd and that 2T/(L1)2T/(L-1) is an integer. Define the block length m:=2T/(L1)m:=2T/(L-1) and partition the horizon into (L1)/2(L-1)/2 disjoint blocks $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\},$ $\forall\ell\in[(L-1)/2].$ Within each block $\ell$, there exists exactly one window of length $l$ whose location is unknown. Fix integers $n,l\geq 1$ with $nl\leq m-2$, and define the candidate window start times (within a block) by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw $\nu=(\nu_{1},\dots,\nu_{(L-1)/2}),$ where $\nu_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}).$ We interpret $\nu_{\ell}$ as the (within-block) start time of the special window in block $\ell$, and define the corresponding absolute start time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each time $t\in\mathcal{I}_{\ell}$, set
@@
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$.
For each window $\tau\in\mathcal{T}$ in block $\ell$, construct the counter and the estimator:
$N_{2}^{(\ell)}(\tau):=\sum_{t=(\ell-1)m+\tau}^{(\ell-1)m+\tau+l-1}\mathds{1}\{a_{t}=2\},\qquad\widehat{\nu}_{\ell}(Z):=\arg\max_{\tau\in\mathcal{T}}N_{2}^{(\ell)}(\tau),$
with ties broken arbitrarily. Define the (blockwise) number of non-satisficing pulls: $M_{\ell}:=\sum_{t\in\mathcal{I}_{\ell}}\mathds{1}\{a_{t}\neq a_{t}^{*}\},$ where $a_{t}^{*}$ is the unique satisficing arm at time $t$. If $M_{\ell}\leq\frac{1}{4}l$, then inside the true window arm $2$ is satisficing for all $l$ rounds, so $N_{2}^{(\ell)}(\nu_{\ell})\ \geq\ l-M_{\ell}\ \geq\ \frac{3}{4}l.$ Consequently, $M_{\ell}\leq\frac{1}{4}l$ implies $\widehat{\nu}_{\ell}(Z)=\nu_{\ell}$, and thus
$\frac{1}{2}\leq\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z)\neq\nu_{\ell}\bigr)\leq\mathbb{P}\bigl(M_{\ell}>\tfrac{1}{4}l\bigr)$ $\leq\frac{4\mathbb{E}[M_{\ell}]}{l}\leq\frac{64\Delta^{2}\mathbb{E}[M_{\ell}]}{\log n}$
Then, $\mathbb{E}[M_{\ell}]\geq\frac{\log n}{128\Delta^{2}}$ for every $\ell\in[(L-1)/2]$. Since each pull of a non-satisficing arm incurs satisficing regret $\Delta$, $\mathbb{E}[R_{T}^{S}(\pi,\nu)]=\Delta\mathbb{E}[M]=\Delta\sum_{\ell\in[(L-1)/2]}\mathbb{E}[M_{\ell}]\geq\frac{(L-1)\log n}{256\Delta}.$
Optimize $n$.
Recall that $nl+2\leq 2T/(L-1)$ and $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$. Ignoring the ceiling, it suffices to choose $n\geq 2$ such that $n\log n\leq 16\Delta^{2}(\frac{2T}{L-1}-2).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}\geq L$, we further have $16\Delta^{2}(\frac{2T}{L-1}-2)>4.$ We will use the following lemma, whose proof is deferred to Section 6.1.1.
@@
6.1.1 Proof of Lemma 1
Let $x:=\lceil\sqrt{y}\rceil$. Then $x\geq\sqrt{y}$, so $\log x\geq\frac{1}{2}\log y$. It remains to show $x\log x\leq y$. Let $n:=\lfloor\sqrt{y}\rfloor$, so $n\geq 2$ and $y\in[n^{2},(n+1)^{2})$, which implies $x=\lceil\sqrt{y}\rceil\leq n+1$. Thus it suffices to prove
@@ 6.1.1 Proof of Lemma 1
For $n=2$, (15) holds since $3\log 3\leq 4$. For $n\geq 3$, define the function $f(u):=u-1-\log(u+1)$ for $u\geq 3$. Then $f^{\prime}(u)=1-\frac{1}{u+1}=\frac{u}{u+1}>0,$ so $f$ is increasing on $3,\infty)$, and $f(n)\geq f(3)=2-\log 4>0$. Hence $\log(n+1)\leq n-1$ for all integers $n\geq 3$, and therefore $(n+1)\log(n+1)\leq(n+1)(n-1)=n^{2}-1\leq n^{2},$ proving (15). Since $y\geq n^{2}$, we conclude $x\log x\leq(n+1)\log(n+1)\leq n^{2}\leq y$. This completes the proof of Lemma 1.

### 25 Apr 2026 v2

@@ 4 Regret Guarantees for Nonstationary Satisficing Bandits
In this section, we establish optimal satisficing regret guarantees for general nonstationary satisficing bandits. Sections 4.1 and 4.2 present the corresponding lower and upper bounds, respectively.
@@ 4.1 Lower Bounds
In the stationary setting, optimal satisficing regret is $T$-independent under realizability (Feng et al. 2024, Rouyer et al. 2024). However, as we show below, such a rate is generally impossible when change points are unknown in advance.
The Hard Instances. Consider a fixed $\Delta>0$, and let $\mathcal{E}_{L,\Delta}$ denote the class of two-armed, piecewise-stationary bandit instances with $L$ segments, where rewards are Gaussian with unit variance and the mean vector on each segment $\ell$ is either $(S+\Delta,S-\Delta)$ or $(S-\Delta,S+\Delta)$, where we recall $S$ is the satisficing threshold. In particular, the identity of the unique satisficing arm changes at every change point. Using the post-interaction reference framework of Section 3.2, we derive a sharp satisficing regret lower bound on $\mathcal{E}_{L,\Delta}$.
@@
When $\Delta$ is a constant and $L\ll\Delta^{2}T$, Theorem 1 implies an $\Omega(L\log T)$ lower bound for general nonstationary satisficing bandits. Theorem 1 reflects the difficulty incurred by unknown change points: minimizing satisficing regret is at least as hard as localizing them, which already incurs a $T$-dependent $\Omega(L\log T)$ cost. The proof, deferred to Section 6.1, is based on Approach I introduced in Section 3.2.
As will become clear in the proof, Theorem 1 requires $L\geq 3$ because the construction (13) relies on a short stationary segment $t_{1},t_{2}]\subseteq[2,T-1]$ with $t_{2}-t_{1}\equiv c$. After the reference is revealed, one can choose two instances that differ only on two short interior intervals, whose total length $2c$ is independent of their locations. This yields the separation condition (14) and hence the reduction (10). When $L=2$, however, there is only one change point. If two instances place it at different times $\tau$ and $\tau^{\prime}$, then they disagree on the entire interval between $\tau$ and $\tau^{\prime}$, whose length grows with $|\tau-\tau^{\prime}|$ (Figure 1). It is therefore unclear how to construct a comparable separated instance family in this case and the possibility of $T$-independent satisficing regret remains.
The next theorem resolves this issue via Approach II, introduced in Section 3.2, and establishes the same scaling for all $L\geq 2$ under a slightly stronger condition relating $\Delta$, $T$, and $L$.
[…]
Figure 1: Schematic illustration: the full horizontal line represents the time horizon $T$, blue/red segments denote the two stationary environments, and the shaded region marks where the two instances disagree.
@@
The full proof of Theorem 2 is deferred to Section 6.2. The theorem also offers a complementary intuition for the $\Omega(L\log T)$ scaling: regardless of the policy, one must acquire $\Omega(\log T)$ information per segment, since the reference can be used to reduce the identification error below a constant. The information-cost bound then implies that acquiring this information necessarily incurs $\Omega(\log T)$ satisficing regret per segment.
Challenge with KL-based Arguments. We also note that it is unclear whether such results can be derived using a KL-based argument. To see the difficulty, consider two instances $\nu,\nu^{\prime}$ that differ only on a disagreement region $D(\nu,\nu^{\prime})\subseteq[T]$. Under Gaussian rewards, the per-round KL divergence at time $t\in D(\nu,\nu^{\prime})$ is of order
$D_{\mathrm{KL}}\!\bigl(\mathcal{N}(\mu_{t}^{\nu}(a),1),\,\mathcal{N}(\mu_{t}^{\nu^{\prime}}(a),1)\bigr)\asymp\bigl(\mu_{t}^{\nu}(a)-\mu_{t}^{\nu^{\prime}}(a)\bigr)^{2}\asymp\Delta^{2},$
so the total transcript KL satisfies $D_{\mathrm{KL}}(P_{\nu}^{\pi},P_{\nu^{\prime}}^{\pi})\approx\Delta^{2}\,|D(\nu,\nu^{\prime})|.$ Thus, the KL divergence is driven primarily by the length of the disagreement region. In particular, it does not directly distinguish whether the samples collected there incur satisficing regret or merely provide information.
Our framework overcomes this difficulty by working instead with (conditional) mutual information, which can be upper bounded by the satisficing regret itself, so that any nontrivial identification task necessarily incurs satisficing regret.
@@ 4.2 Upper Bounds
We now present an algorithm whose satisficing regret matches the lower bounds (up to problem-dependent constants). We begin by defining the windowed statistics used throughout.
Windowed Statistics. For any buffer $\mathcal{B}$ of length $m$, define the dyadic window set $\mathrm{WinSet}(\mathcal{B}):=\{2^{k}:2^{k}\leq m,\ k\in\mathbb{N}\}$. For each $w\in\mathrm{WinSet}(\mathcal{B})$, define the average of the most recent $w$ entries in the buffer $\mathcal{B}$ by $\mathrm{AvgLast}(\mathcal{B},w):=\frac{1}{w}\sum_{s=m-w+1}^{m}\mathcal{B}[s].$ Given a confidence radius $\beta(\cdot)$, define the windowed lower and upper confidence bounds $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B}):=\max_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)-\beta(w)\big\}$ and $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})=\min_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)+\beta(w)\big\}.$
@@
Algorithm 1 maintains a leader state. When $\textsf{leader}=\texttt{None}$, it explores until it certifies that an arm is satisficing. When a leader is present, the algorithm exploits this arm while continuously testing whether it remains satisficing. The key difficulty in piecewise-stationary environments is that, after a change point, stale samples can be misleading. The windowed statistics address this issue: $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})$ takes the most optimistic lower bound over dyadic suffix windows, while $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})$ takes the most pessimistic upper bound. Consequently, if an arm’s current mean exceeds $S$, then eventually $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})\geq S$; conversely, if the current leader becomes non-satisficing, then for a sufficiently recent window we have $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})<S$, triggering a reset and renewed exploration.
We now state the satisficing regret guarantee for Algorithm 1.
@@
Let $\pi$ be Algorithm 1. Under Assumption 1, we have
@@
The $\log T$ factor comes from concentration. With $\beta(w)\asymp\sqrt{\log T/w}$, certifying $\mu>S$ (via $\mathrm{LCB}^{\operatorname{win}}\geq S$) or rejecting $\mu<S$ (via $\mathrm{UCB}^{\operatorname{win}}<S$) requires only $w=\mathcal{O}(\log T/\Delta^{2})$ samples, where $\Delta$ is the relevant gap to the threshold. The windowed bounds also prevent frequent resets: after a reset, if the environment remains unchanged, then with high probability a satisficing leader is not demoted and a non-satisficing candidate is not promoted. Thus, within each stationary segment, the algorithm spends only logarithmically many rounds either finding a satisficing leader or detecting that the current leader has fallen below $S$; afterward it plays a satisficing arm and incurs no further regret. The full proof of Theorem 3 is deferred to Section 6.3.
5 When Is $T$-Independent Satisficing Regret Achievable
It remains unclear which structural conditions permit $T$-independent satisficing regret in nonstationary $K$-armed bandits. A common feature of the hard instances behind the $\Omega(\log T)$ lower bounds in Section 4 is the presence of threshold crossings: an arm may be satisficing early on but later down-cross below $S$ at an unknown time, and the identity of the realizable arm may also switch across segments. Such changes force the learner to repeatedly certify satisficing arms, which incurs a logarithmic cost.
In this section, we identify a regime in which these obstacles disappear. Specifically, if (i) there exists an always-realizable arm that remains strictly satisficing throughout, and (ii) no arm ever transitions from satisficing to non-satisficing (i.e., no down-crossings), then constant satisficing regret—independent of both the horizon $T$ and the number of stationary segments LL—is achievable.
Formally, we impose the following assumptions.
@@ 5 When Is $T$-Independent Satisficing Regret Achievable
Let $\pi$ be Algorithm 2. Under Assumptions 2 and 3, we have
@@
Theorem 4 establishes a constant satisficing regret bound that does not depend on $T$ or $L$, and it recovers the stationary guarantee when $L=1$; see Michel et al. (2022). The key proof idea is that, under Assumptions 2 and 3, the always-realizable arm and any non-satisficing arm can be identified in finite expected time before the latter ever becomes satisficing. The full proof is deferred to Section 6.4. We emphasize that Assumptions 2 and 3 are sufficient (but not necessary) for constant satisficing regret. Characterizing necessary and sufficient structural conditions remains an interesting direction for future work.
@@ 6.1 Proof of Theorem 1
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. Assume $\xi_{t}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume that $L$ is odd and that 2T/(L1)2T/(L-1) is an integer; the even-$L$ case can be handled similarly by adding an initial change point at time $2$. Let $m:=\frac{2T}{L-1}$ and $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\}$ for any $\ell\in[(L-1)/2],$ so the horizon is partitioned into (L1)/2(L-1)/2 disjoint blocks. In each block $\ell$, there is exactly one special window of length $l$, whose start time is unknown. Fix $n,l\geq 1$ with $nl\leq m-2$, and define the candidate within-block start times by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw
$\nu=(\nu_{1},\dots,\nu_{(L-1)/2}),\qquad\nu_{\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}),$
and interpret $\nu_{\ell}$ as the start time of the special window in block $\ell$, with corresponding absolute start time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each $t\in\mathcal{I}_{\ell}$, set
@@
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$ to lower bound partial regret. For each window $\tau\in\mathcal{T}$ in block $\ell$, construct the counter and the estimator:
$N_{2}^{(\ell)}(\tau):=\sum_{t=(\ell-1)m+\tau}^{(\ell-1)m+\tau+l-1}\mathds{1}\{a_{t}=2\},\qquad\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\arg\max_{\tau\in\mathcal{T}}N_{2}^{(\ell)}(\tau),$
with ties broken arbitrarily. Define the blockwise number of non-satisficing pulls: $M_{\ell}:=\sum_{t\in\mathcal{I}_{\ell}}\mathds{1}\{a_{t}\neq a_{t}^{*}\},$ where $a_{t}^{*}$ is the unique satisficing arm at time $t$. Noting that inside the true window arm $2$ is satisficing for all $l$ rounds, if $M_{\ell}\leq\frac{1}{4}l$, then $N_{2}^{(\ell)}(\nu_{\ell})\ \geq\ l-M_{\ell}\ \geq\ \frac{3}{4}l.$ Consequently, $M_{\ell}\leq\frac{1}{4}l$ implies $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})=\nu_{\ell}$, and thus
$\frac{1}{2}\leq\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell}\bigr)\leq\mathbb{P}\bigl(M_{\ell}>\tfrac{1}{4}l\bigr)$ $\leq\frac{4\mathbb{E}[M_{\ell}]}{l}\leq\frac{64\Delta^{2}\mathbb{E}[M_{\ell}]}{\log n}.$
Then, $\mathbb{E}[M_{\ell}]\geq\frac{\log n}{128\Delta^{2}}$ for every $\ell\in[(L-1)/2]$. Since each pull of a non-satisficing arm incurs satisficing regret $\Delta$, the partial regret for block $\ell$ is given by $\Delta\mathbb{E}[M_{\ell}].$ We thus have
$\mathbb{E}[R_{T}^{S}(\pi,\nu)]=\sum_{\ell\in[(L-1)/2]}\Delta\mathbb{E}[M_{\ell}]\geq\frac{(L-1)\log n}{256\Delta}.$
Optimize $n$. Recall that $nl+2\leq 2T/(L-1)$ and $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$. Ignoring the ceiling, it suffices to choose $n\geq 2$ such that $n\log n\leq 16\Delta^{2}(\frac{2T}{L-1}-2).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}\geq L$, we further have $16\Delta^{2}(\frac{2T}{L-1}-2)>4.$ The proof of Theorem 1 is therefore completed by applying Lemma 1.
@@
Proof.
Proof of Lemma 1 Let $x:=\lceil\sqrt{y}\rceil$. Then $x\geq\sqrt{y}$, so $\log x\geq\frac{1}{2}\log y$. It remains to show $x\log x\leq y$. Let $n:=\lfloor\sqrt{y}\rfloor$, so $n\geq 2$ and $y\in[n^{2},(n+1)^{2})$, which implies $x=\lceil\sqrt{y}\rceil\leq n+1$. Thus it suffices to prove
@@
For $n=2$, (15) holds since $3\log 3\leq 4$. For $n\geq 3$, define the function $f(u):=u-1-\log(u+1)$ for $u\geq 3$. Then $f^{\prime}(u)=1-\frac{1}{u+1}=\frac{u}{u+1}>0,$ so $f$ is increasing on $3,\infty)$, and $f(n)\geq f(3)=2-\log 4>0$. Hence $\log(n+1)\leq n-1$ for all integers $n\geq 3$, and therefore $(n+1)\log(n+1)\leq(n+1)(n-1)=n^{2}-1\leq n^{2},$ proving (15). Since $y\geq n^{2}$, we conclude $x\log x\leq(n+1)\log(n+1)\leq n^{2}\leq y$. This completes the proof of Lemma 1. ∎